Circular No. 6953

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

     URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 M4 (LINEAR)

F. B. Zoltowski, Woomera; and P. J. Shelus, McDonald Observatory, have independently reported that another LINEAR discovery, entered on the Minor Planet Center’s NEO Confirmation Page, is a comet. CCD astrometry, that by W. Offutt in response to the Central Bureau’s request:

1998 UT α_{2000} δ_{2000} m_{1} Observer
June 25.34676 19^{h}46^{m}02.^{s}51 −10^{o}04′50′′.6 17.5 Blythe
25.42039 19 45 50.99 −10 05 46.7  _{”}
26.60160 19 42 49.13 −10 20 22.3 14.7 Zoltowski
26.62300 19 42 45.83 −10 20 38.3 14.8  _{”}
26.89961 19 42 03.12 −10 24 08.3 16.4 Casulli
26.91956 19 41 59.94 −10 24 23.6  Tesi
26.92338 19 41 59.32 −10 24 27.0  _{”}
26.92723 19 41 58.78 −10 24 28.4  Casulli
27.02288 19 41 43.99 −10 25 42.0 17.2 Mallia
27.02634 19 41 43.55 −10 25 41.8 16.5  _{”}
27.22651 19 41 12.43 −10 28 14.2 16.6 Offutt
27.24335 19 41 09.77 −10 28 24.6 16.7  _{”}
27.33895 19 40 54.82 −10 29 35.7  Bell
27.34928 19 40 53.14 −10 29 44.6  _{”}
27.36581 19 40 50.60 −10 29 56.6  Shelus
27.37229 19 40 49.57 −10 30 01.5  _{”}
27.59352 19 40 15.38 −10 32 43.2 14.7 Zoltowski
27.61536 19 40 11.94 −10 32 59.8 14.7  _{”}

M. Blythe, F. Shelly and M. Bezpalko (Socorro). 1-m f/2.15 reflector.

Measurers J. Stuart, H. Viggh and R. Sayer. F. B. Zoltowski (Woomera). 0.30-m f/3.3 Schmidt-Cassegrain. Moderately

diffuse coma, faint 1′ tail in p.a. 45^{o} (June 26). V. S. Casulli (Colleverde di Guidonia). 0.40-m f/3.2 reflector. L. Tesi, A. Boattini and M. Tombelli (San Marcello Pistoiese). 0.4-m f/5

reflector. Measurers A. Boattini, G. Forti and L. Tesi. F. Mallia, G. Masi and M. Di Soria (Campocatino). 0.80-m f/4.0 reflector. W. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien. Well-defined conden-

sation; small coma; broad, fanshaped (90^{o} fan) tail with centerline at p.a. 45^{o}, traced to ∼ 30′′ from the nucleus.

G. Bell (Eskridge). 0.45-m f/4.5 reflector. Measurer G. Hug. P. J. Shelus (Fort Davis). 0.76-m f/3.0 reflector. Some nebulosity, hint of

tail extending roughly to the northeast.

1998 June 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden